# Vision Mamba: Efficient Visual Representation Learning with Bidirectional State Space Model

### 17 Jan 2024 v1

@@ 3.5 Efficiency Analysis
Traditional SSM-based methods leverage the fast Fourier transform to boost the convolution operation as shown in Eq. (4). For data-dependent methods, such as Mamba, the SSM operation in Line 11 of Algo. 1 is no longer equivalent to convolution. To address this problem, Mamba and the proposed Vim choose a modern-hardware-friendly way to ensure efficiency. The key idea of this optimization is to avoid the IO-bound and memory-bound of modern hardware accelerators (GPUs).
IO-Efficiency.
The high bandwidth memory (HBM) and SRAM are two important components for GPUs. Among them, SRAM has a larger bandwidth and HBM has a bigger memory size. The standard implementation of Vim’s SSM operation with HBM requires the number of memory IO on the order of $O(\mathtt{B}\mathtt{M}\mathtt{E}\mathtt{N})$. Inspired by Mamba, Vim first reads in $O(\mathtt{B}\mathtt{M}\mathtt{E}+\mathtt{E}\mathtt{N})$ bytes of memory $(\mathbf{\Delta_{o}},\mathbf{{A}_{o}},\mathbf{{B}_{o}},\mathbf{C_{o})}$ from slow HBM to fast SRAM. Then, Vim gets the discrete $\mathbf{\overline{A}_{o}}$, $\mathbf{\overline{B}_{o}}$ of a size of $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$ in SRAM. Last, Vim performs SSM operations in SRAM and writes the output of a size of $(\mathtt{B},\mathtt{M},\mathtt{E})$ back to HBM. This method can help to reduce IOs from $O(\mathtt{B}\mathtt{M}\mathtt{E}\mathtt{N})$ to $O(\mathtt{B}\mathtt{M}\mathtt{E}+\mathtt{E}\mathtt{N})$.
Memory-Efficiency.
To avoid out-of-memory problems and achieve lower memory usage when dealing with long sequences, Vim chooses the same recomputation method as Mamba. For the intermediate states of size $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$ to calculate the gradient, Vim recomputes them at the network backward pass. For intermediate activations such as the output of activation functions and convolution, Vim also recomputes them to optimize the GPU memory requirement, as the activation values take a lot of memory but are fast for recomputation.
Computation-Efficiency.
SSM in Vim block (Line 11 in Algo.1) and self-attention in Transformer both play a key role in providing global context adaptively. Given a visual sequence $\mathbf{T}\in R^{1\times\mathtt{M}\times\mathtt{D}}$ and the default setting $\mathtt{E}=2\mathtt{D}$, the computation complexity of a global self-attention and SSM are:
@@
$\Omega(\text{SSM})=3\mathtt{M}(2\mathtt{D})\mathtt{N}+\mathtt{M}(2\mathtt{D})\mathtt{N}^{2},$ (8)
@@ 4 Experiment
4.1 Image Classification
Settings.
We benchmark Vim on the ImageNet-1K dataset [10], which contains 1.28M training images and 50K validation images from 1,000 categories. All models are trained on the training set, and top-1 accuracy on the validation set is reported. For fair comparisons, our training settings mainly follow DeiT [61]. Specifically, we apply random cropping, random horizontal flipping, label-smoothing regularization, mixup, and random erasing as data augmentations. When training on $224^{2}$ input images, we employ AdamW [44] with a momentum of $0.9$, a total batch size of $1024$, and a weight decay of $0.05$ to optimize models. We train the Vim models for $300$ epochs using a cosine schedule, $1$$\times$$10^{-3}$ initial learning rate, and EMA. During testing, we apply a center crop on the validation set to crop out $224^{2}$ images. Experiments are performed on 8 A800 GPUs.
@@ 4.1 Image Classification
ResNet-18 [25] $224^{2}$ 12M 69.8
ResNet-50 [25] $224^{2}$ 25M 76.2
ResNet-101 [25] $224^{2}$ 45M 77.4
ResNet-152 [25] $224^{2}$ 60M 78.3
ResNeXt50-32x4d [72] $224^{2}$ 25M 77.6
RegNetY-4GF [50] $224^{2}$ 21M 80.0
@@
S4ND-ViT-B [47] $224^{2}$ 89M 80.4
Vim-Ti $224^{2}$ 7M 73.1
Vim-S $224^{2}$ 26M 80.3
Table 1: Comparison with different backbones on ImageNet-1K validation set.
@@
DeepLab v3+ [6] ResNet-101 $512^{2}$ 63M 44.1
UperNet [71] ResNet-50 $512^{2}$ 67M 41.2
UperNet [71] ResNet-101 $512^{2}$ 86M 44.9
UperNet [71] DeiT-Ti $512^{2}$ 11M 39.2
UperNet [71] DeiT-S $512^{2}$ 43M 44.0
UperNet [71] Vim-Ti $512^{2}$ 13M 40.2
UperNet [71] Vim-S $512^{2}$ 46M 44.9
@@
Results.
Tab. 1 compares Vim with ConvNet-based, Transformer-based and SSM-based backbones. Compared to ConvNet-based ResNet [25], Vim demonstrates superior performance. For example, when the parameters are roughly similar, the top-1 accuracy of Vim-Small reaches 80.3, which is 4.1 points higher than that of ResNet50. Compared with the conventional self-attention-based ViT [14], Vim outperforms it by considerable margins in terms of both parameter numbers and classification accuracy. When compared to the highly-optimized ViT-variant, i.e., DeiT [61], Vim surpasses it at different scales with comparable parameter numbers: 0.9 points higher for Vim-Tiny over DeiT-Tiny, and 0.5 points higher for Vim-Small over DeiT-Small. Compared with SSM-based S4ND-ViT-B [47], Vim achieves similar top-1 accuracy with 3$\times$ fewer parameters.
Fig. 1 (b) and (c) compare the FPS and GPU memory of tiny-size Vim and DeiT. Vim demonstrates better efficiency in speed and memory as image resolution grows. Specifically, when the image size is 512, Vim achieves similar FPS and memory as DeiT. As the image size grows to 1248, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory. The pronounced superiority of Vim’s linear scaling in sequence length makes it ready for high-resolution downstream vision applications and long-sequence multi-modality applications.
4.2 Semantic Segmentation
Settings.
We conduct experiments for semantic segmentation on the ADE20K [74] dataset. ADE20K contains 150 fine-grained semantic categories, with 20K, 2K, and 3K images for training, validation, and testing, respectively. We choose UperNet [70] as our base framework. In training, we employ AdamW with a weight decay of $0.01$, and a total batch size of $16$ to optimize models. The employed training schedule uses an initial learning rate of $6$$\times$$10^{-5}$, linear learning rate decay, a linear warmup of $1,500$ iterations, and a total training of $160$K iterations. The data augmentations follow common settings, including random horizontal flipping, random re-scaling within the ratio range $[0.5,2.0]$, and random photometric distortion. During evaluation, we rescale the image to have a shorter side of $512$.
Results.
As shown in Tab. 2, Vim consistently outperforms DeiT across different scales: 1.0 mIoU higher for Vim-Ti over DeiT-Ti, and 0.9 mIoU higher for Vim-S over DeiT-S. Compared to the ResNet-101 backbone, our Vim-S achieves the same segmentation performance with nearly 2$\times$ fewer parameters.
Figure 3: FPS comparison between DeiT-Ti [60] and our Vim-Ti on the commonly used downstream framework. We perform batch inference and benchmark the log-scaled FPS on the architecture with the backbone and FPN. Vim achieves comparable performance to DeiT with a small resolution, i.e., 512. As the input image resolution increases, Vim will have a higher FPS.
@@ 4.3 Object Detection and Instance Segmentation
Settings.
We conduct experiments for object detection and instance segmentation on the COCO 2017 dataset [39]. The COCO 2017 dataset contains 118K images for training, 5K images for validating, and 20K images for testing. We use the canonical Cascade Mask R-CNN [4] as the base framework. For ViT-based backbones, we apply extra configurations (e.g., interleaved window & global attention) to handle the high-resolution images following ViTDet [38]. For SSM-based Vim, we directly use it without any modifications. Other training and evaluation settings are just the same. During training, we employ AdamW with a weight decay of $0.1$, and a total batch size of $64$ to optimize models. The employed training schedule uses an initial learning rate of $1$$\times$$10^{-4}$, linear learning rate decay, and a total training of $380$K iterations. The data augmentations use large-scale jitter data augmentation [19] to 1024$\times$1024 input images. During evaluation, we rescale the image to have a shorter side of 1024.
Results.
@@
We want to highlight that the accuracy superiority is non-trivial since DeiT is equipped with window attention while Vim works in a pure sequence modeling manner. Specifically, to perform representation learning on high-resolution images (i.e., 1024$\times$1024), we follow ViTDet [38] and modify DeiT backbone with the use of 2D window attention, which injects 2D prior and breaks the sequential modeling nature of Transformer. Thanks to the efficiency illustrated in Sec. 3.5, Fig. 1 and Fig. 4, we can directly apply Vim on 1024$\times$1024 input images and learn sequential visual representation for object detection and instance segmentation without need for 2D priors in the backbone.
Figure 4: GPU memory efficiency comparison between DeiT-Ti [60] and our Vim-Ti on the commonly used downstream framework. We perform batch inference and benchmark the GPU memory on the architecture with the backbone and FPN. Vim requires comparable GPU memory to DeiT with a small resolution, i.e., 512. As the input image resolution increases, Vim will use significantly less GPU memory.
@@
Bidirectional SSM + Conv1d 73.1 34.8
Table 4: Ablation study on the bidirectional design. Default setting for Vim is marked in blue.
@@ 4.4 Ablation Study
We ablate the key bidirectional design of Vim, using ImageNet-1K classification and Segmenter [54] on ADE20K semantic segmentation. To fully evaluate the power of learned representation on ImageNet, we use a simple Segmenter head with only 2 layers to perform transfer learning on semantic segmentation. We study these bidirectional strategies:
  • None. We directly adopt Mamba block to process visual sequence with only the forward direction.
@@
  • Bidirectional Block. We pair the stacked blocks. The first block of each pair processes visual sequence in the forward direction, and the second block of each pair processes in the backward direction.
@@
As shown in Tab. 4, directly adopting Mamba block achieves good performance in classification. However, the unnatural unidirectional manner poses challenges in downstream dense prediction. Specifically, the preliminary bidirectional strategy of using Bidirectional Block achieves 7 points lower top-1 accuracy on classification. Yet, it outperforms the vanilla unidirectional Mamba block by 1.3 mIoU on semantic segmentation. By adding extra backward SSM and Conv1d, we achieve similar classification accuracy (73.1 top-1 acc vs. 73.2 top-1 acc) and exceptional segmentation superiority (34.8 mIoU vs. 32.3 mIoU). We use the strategy of Bidirectional SSM + Conv1d as the default setting in our Vim block.

### 10 Feb 2024 v2

@@ 3.5 Efficiency Analysis
Traditional SSM-based methods leverage the fast Fourier transform to boost the convolution operation as shown in Eq. (4). For data-dependent methods, such as Mamba, the SSM operation in Line 11 of Algo. 21 is no longer equivalent to convolution. To address this problem, Mamba and the proposed Vim choose a modern-hardware-friendly way to ensure efficiency. The key idea of this optimization is to avoid the IO-bound and memory-bound of modern hardware accelerators (GPUs).
IO-Efficiency. The high bandwidth memory (HBM) and SRAM are two important components for GPUs. Among them, SRAM has a larger bandwidth and HBM has a bigger memory size. The standard implementation of Vim’s SSM operation with HBM requires the number of memory IO on the order of $O(\mathtt{B}\mathtt{M}\mathtt{E}\mathtt{N})$. Inspired by Mamba, Vim first reads in $O(\mathtt{B}\mathtt{M}\mathtt{E}+\mathtt{E}\mathtt{N})$ bytes of memory $(\mathbf{\Delta_{o}},\mathbf{{A}_{o}},\mathbf{{B}_{o}},\mathbf{C_{o})}$ from slow HBM to fast SRAM. Then, Vim gets the discrete $\mathbf{\overline{A}_{o}}$, $\mathbf{\overline{B}_{o}}$ of a size of $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$ in SRAM. Last, Vim performs SSM operations in SRAM and writes the output of a size of $(\mathtt{B},\mathtt{M},\mathtt{E})$ back to HBM. This method can help to reduce IOs from $O(\mathtt{B}\mathtt{M}\mathtt{E}\mathtt{N})$ to $O(\mathtt{B}\mathtt{M}\mathtt{E}+\mathtt{E}\mathtt{N})$.
Memory-Efficiency. To avoid out-of-memory problems and achieve lower memory usage when dealing with long sequences, Vim chooses the same recomputation method as Mamba. For the intermediate states of size $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$ to calculate the gradient, Vim recomputes them at the network backward pass. For intermediate activations such as the output of activation functions and convolution, Vim also recomputes them to optimize the GPU memory requirement, as the activation values take a lot of memory but are fast for recomputation.
Computation-Efficiency. SSM in Vim block (Line 11 in Algo.21) and self-attention in Transformer both play a key role in providing global context adaptively. Given a visual sequence $\mathbf{T}\in R^{1\times\mathtt{M}\times\mathtt{D}}$ and the default setting $\mathtt{E}=2\mathtt{D}$, the computation complexity of a global self-attention and SSM are:
@@
$\Omega(\text{SSM})=3\mathtt{M}(2\mathtt{D})\mathtt{N}+\mathtt{M}(2\mathtt{D})\mathtt{N},$ (8)
@@ 4 Experiment
ResNet-18 $224^{2}$ 12M 69.8
ResNet-50 $224^{2}$ 25M 76.2
ResNet-101 $224^{2}$ 45M 77.4
ResNet-152 $224^{2}$ 60M 78.3
ResNeXt50-32$\times$4d $224^{2}$ 25M 77.6
RegNetY-4GF $224^{2}$ 21M 80.0
@@
S4ND-ViT-B $224^{2}$ 89M 80.4
Vim-Ti $224^{2}$ 7M 76.1
${}^{\dagger}$ $224^{2}$ 7M 78.3 +2.2
Vim-S $224^{2}$ 26M 80.5
${}^{\dagger}$ $224^{2}$ 26M 81.6 +1.1
Table 1: Comparison with different backbones on ImageNet-1K validation set. ${}^{\dagger}$ represents the model is fine-tuned with our long sequence setting.
4.1 Image Classification
Settings. We benchmark Vim on the ImageNet-1K dataset [9], which contains 1.28M training images and 50K validation images from 1,000 categories. All models are trained on the training set, and top-1 accuracy on the validation set is reported. For fair comparisons, our training settings mainly follow DeiT [60]. Specifically, we apply random cropping, random horizontal flipping, label-smoothing regularization, mixup, and random erasing as data augmentations. When training on $224^{2}$ input images, we employ AdamW [43] with a momentum of $0.9$, a total batch size of $1024$, and a weight decay of $0.05$ to optimize models. We train the Vim models for $300$ epochs using a cosine schedule, 1×1$\times$1 ×$10^{-3}$ initial learning rate, and EMA. During testing, we apply a center crop on the validation set to crop out $224^{2}$ images. Experiments are performed on 8 A800 GPUs.
@@ 4.1 Image Classification
DeepLab v3+ ResNet-101 $512^{2}$ 63M 44.1
UperNet ResNet-50 $512^{2}$ 67M 41.2
UperNet ResNet-101 $512^{2}$ 86M 44.9
UperNet DeiT-Ti $512^{2}$ 11M 39.2
UperNet DeiT-S $512^{2}$ 43M 44.0
UperNet Vim-Ti $512^{2}$ 13M 41.0
UperNet Vim-S $512^{2}$ 46M 44.9
@@
Long Sequence Fine-tuning To make full use of the efficient long sequence modeling power of Vim, we continue to fine-tune Vim with a long sequence setting for 30 epochs after ImageNet pretraining. Specifically, we set a patch extraction stride of $8$ while keeping the patch size unchanged, a constant learning rate of $10^{-5}$, and a weight decay of $10^{-8}$.
Results. Tab. 1 compares Vim with ConvNet-based, Transformer-based and SSM-based backbone networks. Compared to ConvNet-based ResNet [24], Vim demonstrates superior performance. For example, when the parameters are roughly similar, the top-1 accuracy of Vim-Small reaches 80.5, which is 4.3 points higher than that of ResNet50. Compared with the conventional self-attention-based ViT [13], Vim outperforms it by considerable margins in terms of both parameter numbers and classification accuracy. When compared to the highly-optimized ViT-variant, i.e., DeiT [60], Vim surpasses it at different scales with comparable parameter numbers: 3.9 points higher for Vim-Tiny over DeiT-Tiny, and 0.7 points higher for Vim-Small over DeiT-Small. Compared with SSM-based S4ND-ViT-B [46], Vim achieves higher top-1 accuracy with 3$\times$ fewer parameters. After long sequence fine-tuning, ${}^{\dagger}$ and ${}^{\dagger}$ all achieve higher results. Among them, ${}^{\dagger}$ even achieves similar results with DeiT-B. The results demonstrate that Vim can be adapted to longer sequence modeling easily and extract stronger visual representation.
Fig. 1 (b) and (c) compare the FPS and GPU memory of tiny-size Vim and DeiT. Vim demonstrates better efficiency in speed and memory as image resolution grows. Specifically, when the image size is 512$\times$512, Vim achieves similar FPS and memory as DeiT. As the image size grows to 1248$\times$1248, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory. The pronounced superiority of Vim’s linear scaling in sequence length makes it ready for high-resolution downstream vision applications and long-sequence multi-modality applications.
@@
4.2 Semantic Segmentation
Settings. We conduct experiments for semantic segmentation on the ADE20K [73] and use UperNet [70] as the segmentation framework. We provide detailed settings in Sec. B.
Results. As shown in Tab. 2, Vim consistently outperforms DeiT across different scales: 1.8 mIoU higher for Vim-Ti over DeiT-Ti, and 0.9 mIoU higher for Vim-S over DeiT-S. Compared to the ResNet-101 backbone, our Vim-S achieves the same segmentation performance with nearly 2$\times$ fewer parameters.
Figure 3: GPU memory efficiency comparison between DeiT-Ti [59] and our Vim-Ti on the commonly used downstream framework. We perform batch inference and benchmark the GPU memory on the architecture with the backbone and FPN. Vim requires comparable GPU memory to DeiT with a small resolution, i.e., 512$\times$512. As the input image resolution increases, Vim will use significantly less GPU memory.
@@ 4.3 Object Detection and Instance Segmentation
Settings. We conduct experiments for object detection and instance segmentation on the COCO 2017 dataset [38] and use ViTDet [70] as the basic framework. We provide detailed settings in Sec. B.
@@
We highlight that the accuracy superiority is non-trivial since DeiT is equipped with window attention while Vim works in a pure sequence modeling manner. Specifically, to perform representation learning on high-resolution images (i.e., 1024$\times$1024), we follow ViTDet [37] and modify the DeiT backbone with the use of 2D window attention, which injects 2D prior and breaks the sequential modeling nature of Transformer. Thanks to the efficiency illustrated in Sec. 3.5, Fig. 1 and Fig. 3, we can directly apply Vim on 1024$\times$1024 input images and learn sequential visual representation for object detection and instance segmentation without need for 2D priors in the backbone.
Figure 4: FPS comparison between DeiT-Ti [59] and our Vim-Ti on the commonly used downstream framework. We perform batch inference and benchmark the log-scaled FPS on the architecture with the backbone and FPN. Vim achieves comparable performance to DeiT with a small resolution, i.e., 512$\times$512. As the input image resolution increases, Vim has a higher FPS.
@@
Bidirectional SSM + Conv1d 73.9 35.9
Table 4: Ablation study on the bidirectional design. To ensure a fair comparison, we do not use the class token for each experiment. The default setting for Vim is marked in blue.
@@ 4.4 Ablation Study
Bidirectional SSM. We ablate the key bidirectional design of Vim, using ImageNet-1K classification and the Segmenter [53] semantic segmentation framework on ADE20K. To fully evaluate the power of learned representation on ImageNet, we use a simple Segmenter head with only 2 layers to perform transfer learning on semantic segmentation. We study these bidirectional strategies:
  • None. We directly adopt the Mamba block to process visual sequence with only the forward direction.
@@
  • Bidirectional Block. We pair the stacked blocks. The first block of each pair processes visual sequence in the forward direction and the second block of each pair processes in the backward direction.
@@
As shown in Tab. 4, directly adopting the Mamba block achieves good performance in classification. However, the unnatural unidirectional manner poses challenges in downstream dense prediction. Specifically, the preliminary bidirectional strategy of using Bidirectional Block achieves 7 points lower top-1 accuracy on classification. Yet, it outperforms the vanilla unidirectional Mamba block by 1.3 mIoU on semantic segmentation. By adding extra backward SSM and Conv1d, we achieve superior classification accuracy (73.9 top-1 acc vs. 73.2 top-1 acc) and exceptional segmentation superiority (35.9 mIoU vs. 32.3 mIoU). We use the strategy of Bidirectional SSM + Conv1d as the default setting in our Vim block.
Classification Design. We ablate the classification design of Vim, benchmarking on ImageNet-1K classification. We study the following classification strategies:
  • Mean pool. We adopt mean pooling on the output feature from the last Vim block and perform classification on this pooled feature.
  • Max pool. We first adapt the classification head on each token of the visual sequence and then perform max pooling on the sequence to get the classification prediction result.
  • Head class token. Following DeiT [60], we concatenate the class token at the head of the visual sequence and perform classification.
  • Double class token. Based on the head class token strategy, we additionally add a class token at the tail of the visual sequence.
  • Middle class token. We add a class token at the middle of the visual sequence and then perform classification on the final middle class token.
Classification strategy ImageNet top-1 acc.
Mean pool 73.9
Max pool 73.4
Head class token 75.2
Double class token 74.3
Middle class token 76.1
Table 5: Ablation study on the classification design. The default setting for Vim is marked in blue.
As shown in Tab. 5, experiments show that the middle class token strategy can fully exploit the recurrent nature of SSM and the central object prior in ImageNet, demonstrating the best top-1 accuracy of 76.1.
@@ Acknowledgement
Appendix A Visualization
Appendix B Additional Setting
Settings for Semantic Segmentation. We conduct experiments for semantic segmentation on the ADE20K [73] dataset. ADE20K contains 150 fine-grained semantic categories, with 20K, 2K, and 3K images for training, validation, and testing, respectively. We choose UperNet [69] as our base framework. In training, we employ AdamW with a weight decay of $0.01$, and a total batch size of $16$ to optimize models. The employed training schedule uses an initial learning rate of $6$$\times$$10^{-5}$, linear learning rate decay, a linear warmup of $1,500$ iterations, and a total training of $160$K iterations. The data augmentations follow common settings, including random horizontal flipping, random re-scaling within the ratio range $[0.5,2.0]$, and random photometric distortion. During evaluation, we rescale the image to have a shorter side of $512$.
Settings for Object Detection and Instance Segmentation. We conduct experiments for object detection and instance segmentation on the COCO 2017 dataset [38]. The COCO 2017 dataset contains 118K images for training, 5K images for validating, and 20K images for testing. We use the canonical Cascade Mask R-CNN [4] as the base framework. For ViT-based backbones, we apply extra configurations (e.g., interleaved window & global attention) to handle the high-resolution images following ViTDet [37]. For SSM-based Vim, we directly use it without any modifications. Other training and evaluation settings are just the same. During training, we employ AdamW with a weight decay of $0.1$, and a total batch size of $64$ to optimize models. The employed training schedule uses an initial learning rate of $1$$\times$$10^{-4}$, linear learning rate decay, and a total training of $380$K iterations. The data augmentations use large-scale jitter data augmentation [18] to 1024$\times$1024 input images. During evaluation, we rescale the image to have a shorter side of 1024.